# Fixation and extinction in time-fluctuating spatially structured metapopulations

### v1

@@ III.0.1 Competition-dominated dynamics
The coarse-grained competition-dominated dynamics is thus a birth-death process for the $M$-cluster size $i$, with absorbing boundaries at $i=\Omega$ ($M$ fixation) and $i=0$ ($W$ fixation), see Appendix G. In this representation, the $M$ fixation probability in a metapopulation of size $\Omega$, spatially structured as a graph ${\rm G}$, consisting initially of $i$ mutant demes is denoted $\phi_{i}^{{\rm G}}$, and the unconditional (i.e. regardless of whether $M$ or $W$ takes over [7, 67, 66]) mean fixation time (uMFT) denoted $\theta_{i}^{{\rm G}}$. These quantities satisfy the first-step equations [71, 7, 67, 66]
@@
$\theta^{{\rm G}}(m,{\rm G},K)$ $=\frac{1-\gamma}{1-\gamma^{\Omega}}\sum_{k=1}^{\Omega-1}\sum_{n=1}^{k}\frac{\gamma^{k-n}}{T^{+}_{n}(m,{\rm G},K)},$
@@
where $\gamma\equiv T_{i}^{-}/T_{i}^{+}=\rho_{W}/\rho_{M}\approx\exp(-Ks)$ is a quantity independent of $m$ and ${\rm G}$. As noted in Refs. [41, 50, 42] the fixation probability $\phi^{{\rm G}}=\phi$ is therefore independent of the migration rate and spatial structure. This remarkable result stems from the graphs considered here being circulations: There is the same incoming and outgoing migration flow at each deme. In static environments, a generalised circulation theorem ensures that the fixation probability is independent of $m$ and ${\rm G}$ for circulation graphs [47, 41, 42], a feature displayed in the stochastic simulations of Fig. 2(a,c) for the full microscopic model. In excellent agreement with simulation data of Fig. 2(a,c), we find that the $M$ fixation probability increases almost exponentially and approaches $1$ when $Ks\gg 1$, $\phi^{{\rm G}}\approx 1$. This stems from the invasion of $W$ demes being increasingly likely (and the invasion of $M$ demes exponentially less likely) when the average number of migrations ($mK$) increases along with $K$. When $Ks\ll 1$, the competition is effectively neutral, and in this case $\phi^{{\rm G}}\approx 1/\Omega$. In good agreement with simulation results of Fig. 2(b,d), Eq. (12), predicts that the uMFT decreases with the migration rate $\theta^{{\rm G}}\sim 1/m$ and, for given parameters, the uMFT is shortest on cliques, while it is larger on cycles than on grids. Intuitively, for higher $m$ and more connected graphs, migrants spread faster leading to quicker invasion and fixation.
III.0.2 Extinction-dominated dynamics
@@
In the extinction-dominated regime $\psi<1$ with $m\ll 1$, we do not expect any deme invasions in a time $\tau_{E}$. Therefore, the timescale of extinction dynamics is much shorter than that of $M/W$ competition, and site extinction dominates over deme invasion, see Fig. 1e. The dynamics in this extinction-dominated regime is governed by the random extinction of demes, regardless of their type. Following Ref. [59], in the realm of the coarse-grained description outlined in Appendix C, demes are regarded as being either occupied (by either $W$ or $M$ cells) or empty. Deme extinction occurs randomly while empty demes may be recolonised by migrations from occupied neighbouring demes. The coarse-grained dynamics is thus independent of the spatial structure, and can be represented by a birth-death process for the number of occupied demes, assuming that the site extinction and recolonisation occur instantaneously [59], see Appendix C. In this coarse-grained representation, when all demes are initially occupied ($n=K$), the metapopulation mean extinction time (mMET) is given by Eq. (39) in Appendix C. When $\Omega\gg 1$ and $\psi<1$, the leading contribution to the mMET arises from the term $i=n$ in the innermost sum, yielding
@@
where $\gamma_{\text{EM}}\approx 0.577...$ is the Euler-Mascheroni constant. Eq. (13) gives a good approximation of the mMET. As $\psi$ increases with $K$, at the upper-limit of the extinction-dominated limit (where $\psi$ approaches $1$), the mMET grows almost exponentially with $K$ and logarithmically with $\Omega$, $\theta_{E}\approx e^{K}\ln{(\Omega)}/K$. Eq. (39) thus predicts a rapid growth of the mMET when $\psi\gtrsim 1$, as shown in Fig. 2(f,g) where simulation results are found to be in good agreement with Eq. (39).
III.0.3 Intermediate dynamics
In the intermediate regime $\psi\gtrsim 1$ and $m\ll 1$, there is a crossover between the extinction and competition-dominated dynamics. At fixed migration rate $m$, this crossover regime occurs in the range $\ln{(1/m)}\lesssim K\lesssim\ln{(\Omega K/m)}$. Furthermore, the metapopulation is partially occupied for $\ln{(1/m)}\lesssim K\lesssim\ln{(\Omega/m)}$ corresponding to $11<\psi<\Omega-$ 1. These bounds are illustrated by the vertical lines in Fig. 11.
When $\psi\gtrsim 1$, migration and deme extinction occur on the timescale $\tau_{E}$. In the long run, deme recolonisations and extinctions balance each other, yielding a dynamical equilibrium consisting of $\Omega_{\rm occ}=\Omega(1-1/\psi)$ occupied demes and $\Omega-\Omega_{\rm occ}$ empty demes, see Fig. 9. In this regime, the three-state coarse-grained description of the dynamical equilibrium consists of a random mixture of empty demes, and occupied $W$ and $M$ demes, see Fig. 1(d). After a mean time $\theta_{\rm int}^{\rm G}$ the metapopulation reaches the dynamical equilibrium consisting of a fraction $1-\Omega_{\rm occ}/\Omega=1/\psi$ empty demes, and the remaining demes are all of either type $M$ or $W$ with probability $\phi^{G}_{\rm int}$ and $1-\phi_{\rm int}^{\rm G}$, respectively. The dynamical equilibrium is thus defined by the quantities $\psi$, given by Eq. (7), and $\phi_{\rm int}^{\rm G}$ which is the probability that mutants $M$ take over the $\Omega_{\rm occ}=\Omega(1-1/\psi)$ occupied demes in a mean time $\theta_{\rm int}^{\rm G}$. These quantities are derived and discussed in in Appendix F for a metapopulation on an arbitrary regular lattice consisting initially of a single $M$ deme, see (44) and Fig. 11
@@ IV Time-fluctuating environments
Figure 3: (a) Left: single deme in time-switching environment. The carrying capacity $K(t)$ encodes environmental variability by switching between $K=K_{+}$ (mild environment, green / light) and $K(t)=K_{-}<K_{+}$ (harsh environment, orange / dark) at symmetric rate $\nu$ (see also Appendix H). Communities are larger in the mild environment. When $K(t)$ switches at an intermediate rate, each deme experiences bottlenecks at an average frequency $1/(2\nu)$. Right: $n$ and $K$ vs. time in the intermediate switching regime where the size $n$ of a deme undergoes bottlenecks. Parameters are: $K_{+}=200$, $\nu=0.05$ and $K_{-}=100$ (top) and $K_{-}=5$ (bottom). The bottlenecks are weak when $\psi(m,K_{-})\gg 1$ (top, right) where deme extinction is unlikely. When $\psi(m,K_{-})<1$, there are strong bottlenecks and each deme can go extinct in the harsh environment (bottom, right). (b) Clique metapopulation with $\Omega=6$ connected demes (double arrows). All demes have the same time-switching carrying capacity $K(t)$ encoding environmental variability across the metapopulation, with each deme in the same environmental state. (c) Example evolution across two nearest-neighbour demes in a switching environment subject to strong bottlenecks in the intermediate switching regime, see text. Starting in the mild environment where $K=K_{+}$, the carrying capacity switches to $K_{-}$ (harsh environment) after $t\sim 1/\nu$. Following the $K_{+}\to K_{-}$ switch, each deme size decreases and each subpopulation is subject to strong demographic fluctuations and hence prone to extinction. In the absence of recolonisation of empty demes, effective only in the mild state, all demes go extinct. If there is a switch back to the mild environment $K_{-}\to K_{+}$ prior to total extinction, empty demes can be rescued by migration and recolonised by incoming $W$ or $M$ cells from neighbouring demes. In the sketch, an empty deme is recolonised by a mutant in the mild environment and becomes an $M$ deme. The cycle continues until the entire metapopulation consists of only $W$ or $M$ demes, or metapopulation extinction.
@@
Microbial populations generally evolve in time-varying environments, and are often subject to conditions changing suddenly and drastically, e.g. experiencing cycles of harsh and mild environmental states [72, 20, 73, 30, 74, 75, 76, 77, 78, 28, 79, 80, 81, 82, 83], see Fig. 3. These variations cause fluctuations often associated with population bottlenecks, arising when the deme size is drastically reduced, e.g. due to nutrient scarcity or exposure to toxins [77, 28, 29, 27, 31]. Here, environmental variability is encoded in the time-fluctuating carrying capacity $K(t)$ (5) driven by the DMN $\xi(t)\in\{-1,1\}$ [68, 69, 70, 16, 17, 18, 23, 20, 19, 21, 22, 24, 32], see Sec. II. Hence, when $\nu\lesssim 1$, the deme size $n$ tracks $K(t)$ and experiences a bottleneck whenever the carrying capacity switches from $K_{+}$ to $K_{-}$, see Appendix A.2 and Fig. 3(a,right) [16, 17, 18].
@@ IV.0.1 Weak bottlenecks: $\psi(m,K_{-})\gg 1$
Figure 4: Fixation probability $\Phi^{\rm G}$ and mean fixation time $\Theta^{\rm G}$ against switching rate $\nu$ for various parameters. Each panel shows $\Phi^{\rm G}$ vs. $\nu$ (left) and $\Theta^{\rm G}$ vs. $\nu$ (right). Markers show simulation results and lines are predictions of Eq. (24). (a,b) $\Phi^{\rm clique}(\nu)$ and $\Theta^{\rm clique}(\nu)$ for a clique metapopulation and different values of $m$ in (a) and $s$ in (b). (a) $m=10^{-5}$ (red), $m=10^{-4}$ (blue), $m=10^{-3}$ (yellow), and $s=0.01$. (b) $s=10^{-3}$ (red), $s=10^{-2}$ (blue), $s=10^{-1}$ (yellow), and $m=10^{-4}$. Dashed black lines are guides to the eye showing $\Phi_{0,\infty}$ in (a,left) and $\Theta_{0,\infty}$ in (a,right), see text. Other parameters are $\Omega=16$, $K_{+}=200$, and $K_{-}=20$. (c) $\Phi^{\rm G}(\nu)$ and $\Theta^{\rm G}(\nu)$ for clique (red, crosses), cycle (blue, circles), and grid (yellow, triangles). Other parameters are $\Omega=16$, $K_{+}=200$, $K_{-}=20$, $s=0.01$, $m=10^{-4}$. (d) $\Phi^{\rm clique}(\nu)$ and $\Theta^{\rm clique}(\nu)$ for a clique metapopulation with $K_{+}=200$ (red), $K_{+}=500$ (blue), and $K_{+}=1000$ (yellow). Deviations occur for $\Theta$ with $K_{+}=1000$ since the slow-migration condition is not satisfied in the mild environment. Other parameters are $\Omega=16$, $K_{-}=20$, and $s=0.01$, $m=10^{-4}$. In all examples, there is initially a single $M$ deme and $\Omega-1$ others of type $W$, see text.
@@
When $K_{+}>K_{-}$ with $\psi(m,K_{-})\gg 1$, each deme experiences a weak bottleneck at an average frequency $\nu/2$ when $\nu\lesssim 1$, see Fig. 3(a,top). The condition $\psi(m,K_{-})\approx me^{K_{-}}\gg 1$ ensures that deme extinction can be neglected, with metapopulation dynamics dominated by $M/W$ competition. The metapopulation fate can thus be captured by a two-state coarse-grained description similar to that of Sec. III.0.1. Since the deme size in the environment $\xi$, and hence the number of migrating cells, varies with $K(t)$, it is useful to introduce the long-term average deme size in environmental state $\xi=\pm 1$ under switching rate $\nu$ denoted by $\mathcal{N}_{\xi}(\nu)\equiv\mathcal{N}_{\pm}(\nu)$, see below.
@@
We first discuss the metapopulation fate in the limit of slow and fast environmental switching, and then return to the above case of intermediate switching.
@@
When the environment varies very slowly, $\nu\ll 1$, the carrying capacity remains at its initial value, i.e. $K(t)=K_{+}$ or $K(t)=K_{-}$ each with a probability $1/2$, until invasions lead to the fixation of $W$ or $M$. In the slow switching regime, the $M$ fixation probability and uMFT on a metapopulation spatially arranged as a regular graph ${\rm G}$ are respectively denoted by $\Phi_{0}^{\rm G}$ and $\Theta_{0}^{\rm G}$. The quantities are obtained by averaging their static counterparts (12) over the stationary distribution of $K$, yielding for symmetric switching
@@
the harmonic mean of $K_{+}$ and $K_{-}$, with $\mathcal{N}_{\pm}(\infty)\to{\cal K}$ [16, 17, 18, 22], see Appendix A.2. In this fast switching regime, the $M$ fixation probability and uMFT on a metapopulation spatially arranged as a regular graph ${\rm G}$, respectively denoted by $\Phi_{\infty}^{\rm G}$ and $\Theta_{\infty}^{\rm G}$, are obtained by replacing $K$ with ${\cal K}$ in Eq. (12), yielding
@@
From these expressions and Eq. (12), we notice the fixation probability in the regime of slow and fast switching is independent of the migration rate and spatial structure: $\Phi_{0}^{\rm G}=\Phi^{0}$ and $\Phi_{\infty}^{\rm G}=\Phi_{\infty}$. However, the metapopulation uMFT depends explicitly on the migration rate $m$ and ${\rm G}$, with $\Theta_{0}^{\rm G}\sim 1/m$ and $\Theta_{\infty}^{\rm G}\sim 1/m$.
@@
Under intermediate switching rate, when $\nu\lesssim 1$, the coupling of demographic and environmental fluctuations plays a key role, while cell migration depends on the deme size that in turn varies with the environmental state. Here, the metapopulation dynamics cannot be directly related to its static counterpart. The average deme size $\mathcal{N}_{\xi}(\nu)$ depends non-trivially on $\nu$ and $\xi$, and generally needs to be computed by sampling long-time simulations. However, analytical progress can be made by approximating the population size distribution of an isolated deme in the environmental state $\xi$ by the joint probability density $p_{\xi}(\nu,n)$ of the piecewise deterministic Markov process (PDMP) obtained by ignoring demographic fluctuations [84, 85, 16, 17], see Appendix A.2
$p_{\xi}(\nu,n)=\begin{cases}\frac{\mathcal{Z}_{+}}{n^{2}}\left(\frac{K_{+}-n}{n}\right)^{\nu-1}\left(\frac{n-K_{-}}{n}\right)^{\nu}\quad\text{if $\xi=1$,}\\ \frac{\mathcal{Z}_{-}}{n^{2}}\left(\frac{K_{+}-n}{n}\right)^{\nu}\left(\frac{n-K_{-}}{n}\right)^{\nu-1}\quad\text{if $\xi=-1$}.\end{cases}$ (17)

### v2

@@ III.1 Competition-dominated dynamics
These rates, together with the fact that the timescale of intra-deme dynamics is $1/s$ (timescale of fixation of a single isolated deme [64, 7, 65, 66, 67]; see Table 1) can be used to define slow migration in the competition-dominated regime. To this end, we notice that the growth of the cluster of $i$ mutant demes in time $1/s$ is $T^{+}_{i}(m,{\rm G},K)/s=mKE_{{\rm G}}(i)\rho_{M}/(q_{{\rm G}}s)$. [^1] Here, we use the growth of the $M$ cluster to define slow migration as the mutant has a fitness advantage. Therefore, $\rho_{M}>\rho_{W}$ and $T^{+}_{i}(m,\mathrm{G},K)>T_{i}^{-}(m,\mathrm{G},K)$ in general. In the coarse-grained description, slow migration is the regime where invasion can be regarded as being instantaneous, with fixation of a successful $M$ invader occurring before the next invasion. This requires that the average number of successful $M$ invaders in the time for the intra-deme dynamics be less than one, i.e. $T^{+}_{i}(m,{\rm G},K)/s<1$. Therefore, the condition for slow migration here is
$m<\frac{sq_{{\rm G}}}{K\rho_{M}E_{{\rm G}}(i)}\leq\frac{s}{K\rho_{M}},$ (12)
where we have used $q_{\rm G}/E_{\rm G}(i)\leq 1$. When $s\ll 1$ and $Ks\gg 1$, we have $\rho_{M}\sim s$ [64, 7, 65, 66, 67] and there is slow migration when $m\lesssim 1/K$. For typical values used here, e.g. $\Omega=16,K=100$ and $s=0.1$, we estimate that there is slow migration if $m\lesssim 10^{-2}$, which is in line with the values of $m\in[10^{-5},10^{-2}]$ used in our examples, and explains the deviations reported in Fig. 2(d) for larger $m$. The coarse-grained competition-dominated dynamics is thus a birth-death process for the number $i$ of $M$ demes, with absorbing boundaries at $i=\Omega$ ($M$ fixation) and $i=0$ ($W$ fixation); see Appendix G. In this representation, the $M$ fixation probability in a metapopulation of size $\Omega$, spatially structured as a graph $\rm G$}, consisting initially of $i$ mutant demes is denoted $\phi_{i}^{{\rm G}}$, and the unconditional (i.e. regardless of whether $M$ or $W$ takes over [7, 67, 66]) mean fixation time (uMFT) denoted $\theta_{i}^{{\rm G}}$. These quantities satisfy the first-step equations [72, 7, 67, 66]
@@
$\theta^{{\rm G}}(m,K)$ $=\frac{1-\gamma}{1-\gamma^{\Omega}}\sum_{k=1}^{\Omega-1}\sum_{n=1}^{k}\frac{\gamma^{k-n}}{T^{+}_{n}(m,{\rm G},K)},$
@@
where $\gamma\equiv T_{i}^{-}/T_{i}^{+}=\rho_{W}/\rho_{M}\approx\exp(-Ks)$ is a quantity independent of $m$ and $\rm G$}. As noted in Refs. [41, 50, 42] the fixation probability $\phi^{{\rm G}}=\phi$ is therefore independent of the migration rate and spatial structure. This remarkable result stems from the graphs considered here being circulations; see Eq. (5). In static environments, a generalised circulation theorem ensures that the fixation probability is independent of $m$ and $\rm G$} for circulation graphs [47, 41, 42], a feature displayed in the stochastic simulations of Fig. 2(a,c) for the full microscopic model. In excellent agreement with simulation data of Fig. 2(a,c), we find that the $M$ fixation probability increases almost exponentially with $Ks$ and approaches 11 when $Ks\gg 1$, $\phi^{{\rm G}}\approx 1$. This stems from the invasion of $W$ demes being increasingly likely (and the invasion of $M$ demes exponentially less likely) when the average number of migrations ($mK$) increases along with $K$. When $Ks\ll 1$, the competition is effectively neutral, and in this case $\phi^{{\rm G}}\approx 1/\Omega$. In good agreement with simulation results of Fig. 2(b,d), Eq. (14) predicts that the uMFT decreases with the migration rate $\theta^{{\rm G}}\sim 1/m$ and, for given parameters, the uMFT is shortest on cliques, while it is larger on cycles than on grids. Intuitively, for higher $m$ and more connected graphs, migrants spread faster leading to quicker invasion and fixation.
III.2 Extinction-dominated dynamics
@@
In the extinction-dominated regime $\psi<1$ with $m\ll 1$ (slow migration), we do not expect any deme invasions in a time $\tau_{E}$, enabling us to adopt a suitable coarse-grained description. Deme invasions being negligible under slow migration, the timescale of extinction dynamics is much shorter than that of $M/W$ competition, and site extinction dominates over deme invasion; see Fig. 1(e). The dynamics in this extinction-dominated regime is hence governed by the random extinction of demes, regardless of their type. Following Ref. [59], in the realm of the coarse-grained description outlined in Appendix C, demes are regarded as being either occupied (by either $W$ or $M$ cells) or empty. Deme extinction occurs randomly while empty demes may be recolonised by migrations from occupied neighbouring demes, which occur rarely in this regime. Since spatial structure may only influence the dynamics via migrations, the coarse-grained dynamics is largely independent of the spatial structure in this regime where $m\ll 1$, and can be represented by a birth-death process for the number of occupied demes, assuming that the site extinction and recolonisation occur instantaneously [59]; see Appendix C. In this coarse-grained representation, when all demes are initially occupied ($n=K$), the metapopulation mean extinction time (mMET), denoted by $\theta_{E}$, is given by Eq. (39) in Appendix C. When $\Omega\gg 1$ and $\psi<1$, the leading contribution to the mMET arises from the term $i=n$ in the innermost sum, yielding
@@
where $\gamma_{\text{EM}}\approx 0.577$... is the Euler-Mascheroni constant. Eq. (15) gives a good approximation of the mMET as reported in Fig. 8. As $\psi$ increases with $K$, at the upper-limit of the extinction-dominated limit (where $\psi$ approaches $1$), the mMET grows almost exponentially with $K$ and logarithmically with $\Omega$, $\theta_{E}\approx e^{K}\ln{(\Omega)}/K$. Eq. (39) thus predicts a rapid growth of the mMET when $\psi\gtrsim 1$, as shown in Fig. 2(f) where simulation results are found to be in good agreement with Eq. (39).
III.3 Intermediate dynamics
In the intermediate regime $\psi\gtrsim 1$ with $m\ll 1$ (slow migration), there is a crossover between the extinction and competition-dominated dynamics that we can characterise at a coarse-grained level. At fixed migration rate $m$, this crossover regime occurs in the range $\ln{(1/m)}\lesssim K\lesssim\ln{(\Omega K/m)}$; see Appendix F. Furthermore, the metapopulation is partially occupied for $\ln{(1/m)}\lesssim K\lesssim\ln{(\Omega/m)}$ corresponding to $1<\psi<\Omega-1$. These bounds are illustrated by the vertical lines in Fig. 11.
When $\psi\gtrsim 1$, migration and deme extinction occur on the timescale $\tau_{E}$. In the long run, deme recolonisations and extinctions balance each other, yielding a dynamical equilibrium consisting of $\Omega_{\rm occ}=\Omega(1-1/\psi)$ occupied demes and $\Omega-\Omega_{\rm occ}$ empty demes; see Fig. 9. In this regime, the three-state coarse-grained description of the dynamical equilibrium consists of a random mixture of empty demes, and occupied $W$ and $M$ demes; see Fig. 1(d). After a mean time $\theta_{\rm int}^{\rm G}$ the metapopulation reaches the dynamical equilibrium consisting of a fraction $1-\Omega_{\rm occ}/\Omega=1/\psi$ empty demes, and the remaining demes are all of either type $M$ or $W$ with probability $\phi^{G}_{\rm int}$ and $1-\phi_{\rm int}^{\rm G}$, respectively. The dynamical equilibrium is thus defined by the quantities $\psi$, given by Eq. (8), and $\phi_{\rm int}^{\rm G}$ which is the probability that mutants $M$ take over the $\Omega_{\rm occ}=\Omega(1-1/\psi)$ occupied demes, where the unconditional mean fixation time is given by $\theta_{\rm int}^{\rm G}$. These quantities are derived and discussed in Appendix F for a metapopulation on an arbitrary regular lattice consisting initially of a single $M$ deme; see Eq. (44) and Fig. 11.
@@ IV Time-fluctuating environments
Figure 3: (a) Left: single deme in time-switching environment. The carrying capacity $K(t)$ encodes environmental variability by switching between $K=K_{+}$ (mild environment, green / light) and $K(t)=K_{-}<K_{+}$ (harsh environment, orange / dark) at symmetric rate $\nu$ (see also Appendix H). Communities are larger in the mild environment. When $K(t)$ switches at an intermediate rate $\nu\lesssim 1$, each deme experiences bottlenecks prior to deme fixation at an average frequency $\nu/2$; see text. Right: $n$ and $K$ vs. time in the intermediate switching regime where the size $n$ of a deme undergoes bottlenecks. Parameters are: $K_{+}=200$, $\nu=0.05$ and $K_{-}=100$ (top) and $K_{-}=5$ (bottom). The bottlenecks are weak when $\psi(m,K_{-})\gg 1$ (top, right) where deme extinction is unlikely. When $\psi(m,K_{-})<1$, there are strong bottlenecks and each deme can go extinct in the harsh environment (bottom, right). (b) Clique metapopulation with $\Omega=6$ connected demes (double arrows). All demes have the same time-switching carrying capacity $K(t)$ encoding environmental variability across the metapopulation, with each deme in the same environmental state. (c) Example evolution across two nearest-neighbour demes in a switching environment subject to strong bottlenecks in the intermediate switching regime; see text. Starting in the mild environment where $K=K_{+}$, the carrying capacity switches to $K_{-}$ (harsh environment) after $t\sim 1/\nu$. Following the $K_{+}\to K_{-}$ switch, each deme size decreases and each subpopulation is subject to strong demographic fluctuations and hence prone to extinction. In the absence of recolonisation of empty demes, effective only in the mild state, all demes go extinct. If there is a switch back to the mild environment $K_{-}\to K_{+}$ prior to total extinction, empty demes can be rescued by migration and recolonised by incoming $W$ or $M$ cells from neighbouring demes. In the sketch, an empty deme is recolonised by a mutant in the mild environment and becomes an $M$ deme. The cycle continues until the entire metapopulation consists of only $W$ or $M$ demes, or metapopulation extinction.
@@
Microbial populations generally evolve in time-varying environments, and are often subject to conditions changing suddenly and drastically, e.g. experiencing cycles of harsh and mild environmental states [73, 20, 74, 30, 75, 76, 77, 78, 79, 28, 80, 81, 82, 83, 84], see Fig. 3. These variations cause fluctuations often associated with population bottlenecks, arising when the deme size is drastically reduced, e.g. due to nutrient scarcity or exposure to toxins [78, 28, 29, 27, 31]. Here, environmental variability is encoded in the time-fluctuating carrying capacity $K(t)$ of Eq. (6) driven by the DMN $\xi(t)\in\{-1,1\}$ [68, 69, 70, 16, 17, 18, 23, 20, 19, 21, 22, 24, 32]; see Sec. II. Since the dynamics of the deme size $n$ occurs on a timescale of order $1$ (see Appendix A.2), the variable $n$ tracks $K(t)$ and experiences a bottleneck whenever the carrying capacity switches from $K_{+}$ to $K_{-}$ at a rate $\nu\lesssim 1$, see Fig. 3(a,right) [16, 17, 18].
@@ IV.1 Weak bottlenecks: $\psi(m,K_{-})\gg 1$
Figure 4: Fixation probability $\Phi^{\rm G}$ and mean fixation time $\Theta^{\rm G}$ against switching rate $\nu$ for various parameters. Each panel shows $\Phi^{\rm G}$ vs. $\nu$ (left) and $\Theta^{\rm G}$ vs. $\nu$ (right). Markers show simulation results and lines are predictions of Eq. (26). (a,b) $\Phi^{\rm clique}(\nu)$ and $\Theta^{\rm clique}(\nu)$ for a clique metapopulation and different values of $m$ in (a) and $s$ in (b). (a) $m=10^{-5}$ (red), $m=10^{-4}$ (blue), $m=10^{-3}$ (yellow), and $s=0.01$. (b) $s=10^{-3}$ (red), $s=10^{-2}$ (blue), $s=10^{-1}$ (yellow), and $m=10^{-4}$. Dashed black lines are guides to the eye showing $\Phi_{0,\infty}$ in (a,left) and $\Theta_{0,\infty}$ in (a,right); see text. Other parameters are $\Omega=16$, $K_{+}=200$, and $K_{-}=20$. (c) $\Phi^{\rm G}(\nu)$ and $\Theta^{\rm G}(\nu)$ for clique (red, crosses), cycle (blue, circles), and grid (yellow, triangles) metapopulations. Other parameters are $\Omega=16$, $K_{+}=200$, $K_{-}=20$, $s=0.01$, $m=10^{-4}$. (d) $\Phi^{\rm clique}(\nu)$ and $\Theta^{\rm clique}(\nu)$ for a clique metapopulation with $K_{+}=200$ (red), $K_{+}=500$ (blue), and $K_{+}=1000$ (yellow). Deviations occur for $\Theta$ with $K_{+}=1000$ since the slow-migration condition is not satisfied in the mild environment. Other parameters are $\Omega=16$, $K_{-}=20$, and $s=0.01$, $m=10^{-4}$. In all examples, there is initially a single $M$ deme and $\Omega-1$ others of type $W$; see text.
@@
When $K_{+}>K_{-}$ with $\psi(m,K_{-})\gg 1$, each deme experiences a weak bottleneck at an average frequency $\nu/2$ when $\nu\lesssim 1$ [16, 17, 18]; see Fig. 3(a,top). The condition $\psi(m,K_{-})\approx me^{K_{-}}\gg 1$ ensures that deme extinction can be neglected, with metapopulation dynamics dominated by $M/W$ competition. The metapopulation fate can thus be captured by a two-state coarse-grained description similar to that of Sec. III.1. Since the deme size in the environment $\xi$, and hence the number of migrating cells, varies with $K(t)$, it is useful to introduce the long-term average deme size in environmental state $\xi=\pm 1$ under switching rate $\nu$ denoted by $\mathcal{N}_{\xi}(\nu)\equiv\mathcal{N}_{\pm}(\nu)$; see below.
@@
We first discuss the metapopulation fate in the limit of slow and fast environmental switching, and then return to the above case of weak bottlenecks with $\nu\lesssim 1$.
@@
When the environment varies very slowly, $\nu\ll 1$, the carrying capacity remains at its initial value, i.e. $K(t)=K_{+}$ or $K(t)=K_{-}$ each with a probability $1/2$, until invasions lead to the fixation of $W$ or $M$. In other words, the time between switches is longer than the mean fixation time in each environment such that $\nu\max(\theta^{\rm G}(m,K_{+}),\theta^{\rm G}(m,K_{-}))<1$, where $\theta^{\rm G}(m,K)$ is given by Eq. (14). In the slow switching regime, the $M$ fixation probability and uMFT on a metapopulation spatially arranged as a regular graph $\rm G$} are respectively denoted by $\Phi_{0}^{\rm G}$ and $\Theta_{0}^{\rm G}$. The quantities are obtained by averaging their static counterparts (14) over the stationary distribution of $K$, yielding for symmetric switching
@@
the harmonic mean of $K_{+}$ and $K_{-}$, with ${\cal K}\approx 2K_{-}$ if $K_{-}\gg 1$ and $\mathcal{N}_{\pm}(\infty)\to{\cal K}$ when ${\cal K}\gg 1$  [16, 17, 18, 22]; see Appendix A.2. In this fast switching regime, the $M$ fixation probability and uMFT on a metapopulation spatially arranged as a regular graph $\rm G$}, respectively denoted by $\Phi_{\infty}^{\rm G}$ and $\Theta_{\infty}^{\rm G}$, are obtained by replacing $K$ with $\cal K$} in Eq. (14), yielding
@@
From these expressions and Eq. (14), we notice the fixation probability in the regime of slow and fast switching is independent of the migration rate and spatial structure: $\Phi_{0}^{\rm G}=\Phi_{0}$ and $\Phi_{\infty}^{\rm G}=\Phi_{\infty}$. However, the metapopulation uMFT depends explicitly on the migration rate $m$ and $\rm G$}, with $\Theta_{0}^{\rm G}\sim 1/m$ and $\Theta_{\infty}^{\rm G}\sim 1/m$.
@@
Under intermediate switching rate, when $\nu\lesssim 1$, the coupling of demographic and environmental fluctuations plays a key role, while cell migration depends on the deme size that in turn varies with the environmental state. Here, the metapopulation dynamics cannot be directly related to its static counterpart. The average deme size $\mathcal{N}_{\xi}(\nu)$ depends non-trivially on $\nu$ and $\xi$, and generally needs to be computed by sampling long-time simulations. However, analytical progress can be made by approximating the distribution of the size $n$ of an isolated deme in the environmental state $\xi$ by the joint probability density $p_{\xi}(n;\nu)$ of the piecewise deterministic Markov process (PDMP), where $n$ and $\xi$ are variables, and $\nu$ a parameter, obtained by ignoring demographic fluctuations [85, 86, 16, 17]; see Appendix A.2:
[…]